1. Answer with octyl ethanoate as the esterfication reaction (alcohol: 1-octanol, 3.25ml and carboxylic acid: ethanoic acid, 1.25ml)

(a) Write a complete chemical equation for the esterfication reaction, use structural formulas and write the name of each compound under the formula. (b) define if the esther is polar or non-polar (c) determine the molar ratio of the two reactants, alcohol and carboxylic acid using the balanced chemical reaction (d) calculate the molar mass of the alcohol and carboxylic acid (e) calculate the mass of alcohol and carboxylic acid using density of 0.8276g/ml for 1-octanol and 1.0492g/ml for ethanoic acid (f) calculate moles of each reactant used in the reaction (g) compare theoretical molar ratios to actual values used in the reaction (h) draw structure for propyl ethanoate (i) name and draw the reactants that make up the compound
Answer #1

Ans 1a:

Reaction: 1 1-octanol + 1 ethanoic acid ⇋ 1 octylethanoate + 1 H2O

Ans 1b:

Because of the presence of ‘-COO-‘ group esters are polar molecules. They do not have ‘H’ capable forming hydrogen bond. But, because of the presence ‘O’, they can act as hydrogen bond acceptors. This explains its water solubility to certain extent.

Esters are less polar than alcohols and more polar than ethers.

Ans 1c:

It can be seen from the above reaction that alcohol and carboxylic acids are needed in 1:1 ratio for esterification reaction (theoretically).

Ans 1d:

Molar mass 1-octanol (C8H18O) = 130.23 g/mol

Molar mass of ethanoic acid (CH3COOH) = 60.05 g/mol

Ans 1e:

Density = mass/volume

Mass = Density x volume

Mass of 1-octanol taken = 1.0495 g/mL x 3.25 mL = 3.411 g

Mass of ethanoic acid taken = 0.8276 g/mL x 1.25 mL = 1.035 g

Ans 1f:

Moles = mass/molar mass

Moles of 1-octanol taken = 3.411/130.23 = 0.0262 mol

Moles of ethanoic acid taken = 1.035/60.05 = 0.0172 mol

Ans 1g:

Theoretical molar ratio

Actual molar ratio

1-octanol : ethanoic acid

1:1

1.52 : 1

Conclusion: Alcohol is taken in excess in the actual reaction. OR, ethanoic acid is taken less. So, ethanoic acid is a limiting reagent.

Ans 1h:

Ans 1i: